Since the activity performed in the afternoon about the realignment of the INJ SLED (59170) needs to be completed, we decided to start the HWS-DET compensation automatic measurement on NI CP.
After some preliminary tests performed by Diego, at 19.35 UTC the automatic HWS-DET compensation measurement started.
We checked the firsts actions performed by the automation:
-
19.35 UTC all the CO2 actuators were switched OFF to cool down the ITMs before the measurement (Fig. 1)
-
21.34 UTC (2 h after) the HWS DET acquisition was started (folder 20230307T2234 created; num ref img = 100, num live img = 100, time sleep = 80 sec, num live loops = 200)
-
21.44 UTC CHs ON (Fig. 1 for the opening of CHs flip mirrors and Fig. 2 for the 1° and the 20th WF, in which NI CH curvature start to appear)
In Fig. 1, the reduction of the power visible in channel “TCS_WI_CO2_POWER_CH_PICKOFF_mean” immediately after the flip mirrors were open, was an action performed manually, to restore the current value of WI CH (= 0.065 W), that was previously raised up to 0.165 mW to ease the HWS-INJ SLED beam re-alignment attempt. It will not affect the HWS-DET compensation measurement since it involves the WI TM.
The automatic HWS-DET compensation measurement will continue during the night. The data will be checked and analyzed in the morning.